### v2

# Value Improved Actor Critic Algorithms

Abstract
To learn approximately optimal acting policies for decision problems, modern Actor Critic algorithms rely on deep Neural Networks (DNNs) to parameterize the acting policy and greedification operators to iteratively improve it. The reliance on DNNs suggests an improvement that is gradient based, which is per step much less greedy than the improvement possible by greedier operators such as the greedy update used by Q-learning algorithms. On the other hand, slow and steady changes to the policy can also be beneficial for the stability of the learning process, resulting in a tradeoff between greedification and stability. To address this tradeoff, we propose to extend the standard framework of actor critic algorithms with value-improvement: a second greedification operator applied only when updating the policy’s value estimate. In this framework the agent can evaluate non-parameterized policies and perform much greedier updates while maintaining the steady gradient-based improvement to the parameterized acting policy. We prove that this approach converges in the popular analysis scheme of Generalized Policy Iteration in the finite-horizon domain. Empirically, incorporating value-improvement into the popular off-policy actor-critic algorithms TD3 and SAC significantly improves or matches performance over their respective baselines, across different environments from the DeepMind continuous control domain, with negligible compute and implementation cost.
1 Introduction
The objective of Reinforcement Learning (RL) is to learn acting policies that, when executed, maximize the expected return (i.e. value) in a given task. Modern RL methods of the Actor-Critic (AC) family (e.g. Schulman et al., 2017; Fujimoto et al., 2018; Haarnoja et al., 2018b; Abdolmaleki et al., 2018) use deep neural networks to parameterize the acting policy, which is iteratively improved using variations of stochastic gradient-descent (SGD) based policy improvement operators such as the policy gradient (Sutton et al., 1999). These operators rely on greedification: modifying the policy such that it maximizes the current evaluation (Sutton & Barto, 2018). In gradient-based optimization the magnitude of the update to the policy - the level of greedification - is governed by the learning rate, which cannot be tuned independently to induce the maximum greedification possible at every step, the greedy update $\pi(s)=\operatorname*{arg\,max}_{a}Q^{\pi}(s,a)$. Similarly, executing $N$ repeating gradient steps with respect to the same batch will encourage the parameters to over-fit to the batch (as well as being computationally intensive) and is thus not a feasible alternative. For these reasons, the greedification of DNN-based policies is typically slow compared to, for instance, the $\operatorname*{arg\,max}$ greedification used in Policy Iteration (Sutton & Barto, 2018) and Q-learning (Mnih et al., 2013). While limited greedification can slow down learning, previous work has shown that too much greedification can cause instability in the learning process through overestimation bias (see van Hasselt et al., 2016; Fujimoto et al., 2018), which can be addressed through softer, less-greedy updates (Fox et al., 2016). This leads to a direct tradeoff between greedification and learning stability.
Previous work partially addresses this tradeoff by decoupling the policy improvement into two steps. First, an improved policy with controllable greediness is explicitly produced by a greedification operator as a target. Second, the acting policy is regressed against this target using supervised learning loss, such as cross-entropy. This target policy is usually not a DNN, and can be for instance a Monte Carlo Tree Search-based policy, a variational parametric distribution, or a nonparametric model (see Haarnoja et al., 2018b; Abdolmaleki et al., 2018; Grill et al., 2020; Hessel et al., 2021; Danihelka et al., 2022). Unfortunately, this approach does not address the tradeoff fully. The parameterized acting policy is still improved with gradient-based optimization which imposes similar limitations on the rate of change to the acting policy.
To better address this tradeoff, we propose to decouple the acting policy from the evaluated policy (the policy evaluated by the critic), and apply greedification independently to both. This allows for (i) the evaluation of policies that need not be parameterized and can be arbitrarily greedy, while (ii) maintaining the slower policy improvement to the acting policy that is suitable for DNNs and facilitates learning stability. We refer to an update step which evaluates an independently-improved policy as a value improvement step and to the approach as Value-Improved Actor Critic (VIAC). Since this framework diverges from the assumption made by the majority of RL methods (evaluated policy $\equiv$ acting policy) it is unclear whether this approach converges and for which greedification operators. Our first result is that no standard RL approach can guarantee convergence for all greedification operators (and by extension improvement operators) because the definition of policy improvement allows for infinitesimal improvement. To classify operators that guarantee convergence, we identify necessary and sufficient conditions for greedification to guarantee convergence of a family of Generalized Policy Iteration algorithms (Sutton & Barto (2018), sometimes called specifically Modified or Optimistic Policy Iteration (Tsitsiklis, 2002; Smirnova & Dohmatob, 2019), a popular setup for underlying-convergence analysis of AC algorithms).
We prove convergence for this class of operators and this class of Generalized and Value-Improved Generalized Policy Iteration algorithms in finite-horizon MDPs. Prior work has shown that this setup converges for specific operators, as well as for all operators that induce deterministic policies (see Williams & Baird III, 1993; Tsitsiklis, 2002; Bertsekas, 2011; Smirnova & Dohmatob, 2019). Our result complements prior work by extending convergence to stochastic policies and a large class of practical operators, such as the operator developed for the Gumbel MuZero algorithm (Danihelka et al., 2022). We demonstrate that incorporating value-improvement into practical algorithms can be beneficial with experiments in Deep Mind’s control suite (Tassa et al., 2018) with the popular off-policy AC algorithms TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018b), where in all environments tested VI-TD3/SAC significantly outperform or match their respective baselines.
2 Background
The reinforcement learning problem is formulated as an agent interacting with a Markov Decision Process (MDP) $\mathcal{M}(\mathcal{S},\mathcal{A},P,R,\rho,H)$, where $\mathcal{S}$ a discrete state space, $\mathcal{A}$ a discrete action space, $P:\mathcal{S}\times\mathcal{A}\to\mathscr{P}(\mathcal{S})$ is a conditional probability measure over the state space that defines the transition probability $P(s,a)$. The immediate reward $R(s,a)$ is a state-action dependent bounded random variable. Initial states are sampled from the start-state distribution $\rho$. In finite horizon MDPs, $H$ specifies the length of a trajectory in the environment. Many RL setups and algorithms consider the infinite horizion case, where $H\to\infty$, but for simplicity’s sake our theoretical analysis in Section 3 remains restricted to finite horizons. The objective of the agent is to find a policy $\pi:\mathcal{S}\to\mathscr{P}(\mathcal{A})$, a distribution over actions at each state, that maximizes the objective $J$, the expected return from the starting state distribution $\rho$. We denote the set of all possible policies with $\Pi$. This quantity can also be written as the expected state value $V^{\pi}$ with respect to starting states $s_{0}$:
$J(\pi)=\mathbb{E}\big{[}V^{\pi}(s_{0})\,\big{|}\!\!\begin{array}[]{c}\scriptstyle s_{0}\sim\rho\end{array}\!\!\big{]}=\mathbb{E}\Big{[}{\textstyle\sum\limits_{t=0}^{H-1}}\gamma^{t}r_{t}\,\Big{|}\!\!\begin{array}[]{c}\scriptstyle s_{0}\sim\rho,\,s_{t+1}\sim P(s_{t},a_{t})\\[-2.84526pt] \scriptstyle a_{t}\sim\pi(s_{t}),\,r_{t}\sim R(s_{t},a_{t})\end{array}\!\!\Big{]}\,.$
The discount factor $0<\gamma\leq 1$ is traditionally set to 1 in finite horizon MDPs. The state value $V^{\pi}$ can also be used to define a state-action $Q$-value and vice versa, i.e. $\forall s\in\mathcal{S},\forall a\in\mathcal{A}$:
$Q^{\pi}(s,a)=\mathbb{E}\Big{[}r+\gamma V^{\pi}(s^{\prime})\,\Big{|}\,\!\!\begin{array}[]{c}\scriptstyle r\sim R(s,a)\\[-2.84526pt] \scriptstyle s^{\prime}\sim P(s,a)\end{array}\!\!\Big{]}\,,\qquad V^{\pi}(s)=\mathbb{E}\big{[}Q^{\pi}(s,a)\,\big{|}\!\!\begin{array}[]{c}\scriptstyle a\sim\pi(s)\end{array}\!\!\big{]}\,.$
Policy Improvement
Directly finding the policy that maximizes the objective $\operatorname*{arg\,max}_{\pi}J(\pi)$ is generally intractable. For that reason, the majority of RL and Dynamic Programming (DP) approaches rely on the iteration of two interleaved processes: first, a policy $\pi$ (an actor) is evaluated with exact $Q^{\pi}$ or approximate $q\approx Q^{\pi}$ (the critic). Second, the policy is improved using $Q^{\pi}$. In RL, this approach is generally referred to as an Actor-Critic (AC), mirrored by the DP approach of (generalized) Policy Iteration. We include a concrete variation of a Generalized Policy Iteration algorithm (sometimes called Optimistic or Modified Policy Iteration, see Bertsekas (2011)) in Algorithm 1.
Algorithm 1 Generalized Policy Iteration
1:For starting functions $q\in\mathcal{Q},\,\pi\in\Pi$ greedification operator $\mathcal{I}$, $k\geq 1$ and $\epsilon>0$
2:while $|\sum_{a\in\mathcal{A}}\big{(}\pi(a|s)q(s,a)\big{)}-\max_{b}q(s,b)|>\epsilon$ or $|q(s,a)-\mathcal{T}^{*}q(s,a)|>\epsilon,\forall s\in\mathcal{S}$ do
3:     $\pi(s)\leftarrow\mathcal{I}(\pi,q)(s),\,\forall s\in\mathcal{S}$
4:     $q(s,a)\leftarrow(\mathcal{T}^{\pi})^{k}q,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
The step $q(s,a)\leftarrow(\mathcal{T}^{\pi})^{k}q,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ denotes $k$ repeating Bellman updates $q_{i+1}(s,a)=\mathcal{T}^{\pi}q_{i}=\mathbb{E}[R(s,a)]+\gamma\mathbb{E}_{s^{\prime}\sim P}[\sum_{a^{\prime}\in\mathcal{A}}\pi(a^{\prime}|s^{\prime})q(s^{\prime},a^{\prime})],i=1,\dots,k$. When $k>H$ the evaluation is exact and the algorithm reduces to Policy Iteration. The set $q\in\mathcal{Q}$ denotes all bounded functions $q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. To drive the iterative improvement of the policy (line 3 in Algorithm 1) these approaches rely on the policy improvement Theorem (Sutton & Barto, 2018):
Theorem 1 (Policy Improvement).
Let $\pi$ and $\pi^{\prime}$ be two policies such that $\forall s\in\mathcal{S}$:
$\sum_{a\in\mathcal{A}}Q^{\pi}(s,a)\pi^{\prime}(a|s)$ $\geq\sum_{a\in\mathcal{A}}Q^{\pi}(s,a)\pi(a|s):=V^{\pi}(s).$ (1)
[…]
In addition, if there is strict inequality of Equation 1 at any state, then there must be strict inequality of Equation 2 at at least one state.
See Sutton & Barto (2018) for proof. The policy improvement theorem connects the policy improvement property (Equation 2) with an easier to optimize objective (Equation 1) which locally for states $s$ searches for a policy $\pi^{\prime}(s)$ that takes better actions with respect to the value of $\pi$. We refer to the process of searching for such a policy as greedification, call $\pi^{\prime}$ greedier than $\pi$ if the inequality holds, and refer to the policy that maximizes the inequality as the greedy policy.
Greedification operators
In RL, operators that produce greedier policies $\pi^{\prime}$ are often referred to as greedification operators (Chan et al., 2022) or more generally as policy improvement operators (Li et al., 2023). To distinguish between operators that generally produce policy improvement (inequality 2) and operators that specifically rely on greedification (inequality 1) for policy improvement, we explicitly distinguish between greedification operators and policy improvement operators:
Definition 1 (Greedification Operators).
If an operator $\mathcal{I}:\Pi\times\mathcal{Q}\to\Pi$ satisfies:
$\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)\geq\sum_{a\in\mathcal{A}}\pi(a|s)q(s,a),\quad\forall\pi\in\Pi,\,\forall q\in\mathcal{Q},\,\forall s\in S,$ (3)
as well as $\exists s\in\mathcal{S}$ such that:
$\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)>\sum_{a\in\mathcal{A}}\pi(a|s)q(s,a),\quad\forall\pi\in\Pi,\,\forall q\in\mathcal{Q},$ (4)
unless the policy is already an $\operatorname*{arg\,max}$ policy, we call it a greedification operator.
Respectively, policy improvement operators are more generally operators that satisfy the inequality in Equation 2. Note that every greedification operator is a policy improvement operator when $q=Q^{\pi}$. However the inverse is not true - not every policy improvement operator is a greedification operator (random mutation operators for example can produce policy improvement without producing greedification, see Supplementary Materials C.2 for a concrete example).
Perhaps the most famous greedification operator is the greedy operator $\mathcal{I}_{\operatorname*{arg\,max}}(\pi,q)(s)=\operatorname*{arg\,max}_{a}q(s,a)$, which drives foundational algorithms such as Value Iteration, Policy Iteration and Q-learning (Sutton & Barto, 2018). The majority of actor critic algorithms on the other hand rely on variations of the policy gradient operator (Sutton et al., 1999), which is well suited for the greedification of parameterized policies.
Deterministic operators
While the policy gradient generally relies on stochastic policies, Silver et al. (2014) propose a deterministic variation of the gradient of the stochastic policy. This deterministic policy gradient is used by popular off-policy AC algorithms such as TD3 (Fujimoto et al., 2018) and TD7 (Fujimoto et al., 2023). In order to better understand the convergence properties of AC algorithms, Williams & Baird III (1993) investigate the convergence properties of Generalized Policy Iteration algorithms that are batch-update based (asynchronous), with all greedification operators that produce policies that are deterministic, $\mathcal{I}_{det}$.
Sampling-based operators
The majority of AC algorithms remain reliant on stochastic policies however, with many operators approximating a greedier policy from samples. Sampling based policies have certain advantages over parameterized policies: they can represent arbitrarily complex probability distribution, given sufficient samples, and they need not rely on SGD for greedification. On the other hand, sampling based policies are harder to maintain between iterations compared to parameterized policies. A rudimentary example is an operator that approximates the $\operatorname*{arg\,max}$ by sampling $N$ actions from a policy $\pi$ and outputting the maximizing action $\operatorname*{arg\,max}_{i\leq N}q(s,a_{i})$. MPO (Abdolmaleki et al., 2018) uses a greedification operator which samples actions from a parameterized policy and reweighs them using the critic to produce greedification. The Gumbel MuZero algorithm (Danihelka et al., 2022) uses the greedification operator $\mathcal{I}_{gmz}(\pi,q)(s)=\mathrm{softmax}(\sigma(q(s,\cdot))+\log\pi(s))$ (for $\sigma$ a monotonically increasing transformation).
Implicit operators
The above mentioned operators all produce an explicit greedier policy $\pi^{\prime}=\mathcal{I}(\pi,q)$. Recently, Kostrikov et al. (2022) proposed that it is also possible to produce implicit greedification, by training a critic to approximate the value of a greedier policy without that policy being explicitly defined. They demonstrate that by training a critic $v_{\psi}$ with the asymmetric expectile loss $\mathcal{L}_{2}^{\tau}$ on a data set $\mathcal{D}$ drawn with some arbitray sampling policy $\pi$,
$\mathcal{L(\theta)}=\mathbb{E}\big{[}\mathcal{L}_{2}^{\tau}\big{(}v_{\psi}(s),Q^{\pi}(s,a)\big{)}\big{|}s,a\sim\mathcal{D}\big{]}\,,\quad\mathcal{L}_{2}^{\tau}(x,y)=|\tau-\mathbbm{1}_{y-x<0}|\,(y-x)^{2}\,,$ (5)
for $\tau>\frac{1}{2}$ the critic $v_{\psi}(s)$ directly estimates the value of a policy than is greedier than $\pi$, with $\tau\to 1$ corresponding to the value of an $\operatorname*{arg\,max}$ policy. This operator is then used to drive their Implicit Q-learning (IQL) algorithm for offline-RL, where the $\mathcal{L}_{2}^{\tau}$ enables the critic to approximate the value of an optimal policy without the bootstrapping of actions that are out of the training distribution.
3 Value Improved Generalized Policy Iteration Algorithms
For theoretical convergence analysis, we begin by formulating a novel DP framework that decouples the improvement of the acting policy from that of the evaluated policy in Algorithm 2, which we call Value-Improved Generalized Policy Iteration. Modifications to the original algorithm in blue.
Algorithm 2 Value-Improved Generalized Policy Iteration
1:For starting vectors $q\in\mathcal{Q},\,\pi\in\Pi$, policy improvement operators $\mathcal{I}_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{I}_{2}}$, $k\geq 1$
2:while $|\sum_{a\in\mathcal{A}}\big{(}\pi(a|s)q(s,a)\big{)}-\max_{b}q(s,b)|>0,\forall s\in\mathcal{S}$ and $|q(s,a)-\mathcal{T}^{*}q(s,a)|>0$ do
3:     $\pi(s)\leftarrow\mathcal{I}_{1}(\pi,q)(s),\,\forall s\in\mathcal{S}$
4:     $q(s,a)\leftarrow(\mathcal{T}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{I}_{2}(\pi,q)}})^{k}q,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
Since the acting and evaluated policies are improved with different operators $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, it is not apparent whether $\pi$ of Algorithm 2 converges to the optimal policy, i.e. is decoupling the policies sound. Therefore, we must first establish for which pairs of operators this process converges. A fundamental result in RL is that policy iteration algorithms converge for any policy improvement operator (and by extension, greedification operator) that produces deterministic policies. This holds because a finite MDP has only a finite number of deterministic policies through which the policy iteration process iterates (Sutton & Barto, 2018). This result however does not generalize to operators that produce stochastic policies, which are used by many practical RL algorithms such as PPO (Schulman et al., 2017), MPO (Abdolmaleki et al., 2018), SAC (Haarnoja et al., 2018b), and Gumbel MuZero (Danihelka et al., 2022).
Theorem 2 (Improvement is not enough).
Policy improvement is not a sufficient condition for the convergence of Policy Iteration algorithms (Algorithm 1 with exact evaluation) to the optimal policy for all starting policies $\pi_{0}\in\Pi$ in all finite-state MDPs.
Proof sketch.
With stochastic policies, an infinitesimal policy improvement is possible, which can satisfy the policy improvement condition at every step and yet converge in the limit to policies that are not $\operatorname*{arg\,max}$ policies. Since every optimal policy is an $\operatorname*{arg\,max}$ policy, Policy Iteration with such operators cannot be guaranteed to converge to the optimal policy. For a complete proof see Supplementary Materials C.1.
Why is this a problem?
Many algorithms are motivated by greedification, but this is not sufficient to establish that the resulting policy improvement will lead to an optimal policy. For that reason, convergence for these algorithms must generally proven individually for each new operator (e.g., see MPO and GreedyAC Chan et al. (2022)), which is often an arduous and nontrivial process.
Furthermore, Theorem 2 and its underlying intuition highlight a critical gap: we currently lack guiding principles for designing novel greedification operators in the form of necessary and sufficient conditions for convergence to the optimal policy. To illustrate that this can lead to problems in practice, we show in Supplementary Materials C.5 and C.6 that viable choices of the transformation $\sigma$ used by the Gumbel MuZero operator $\mathcal{I}_{gmz}$ can render this operator sufficient or insufficient.
To address this problem, we identify a necessary condition and two independent sufficient conditions for greedification operators, such that they induce convergence of Algorithm 1.
Definition 2 (Necessary Greedification).
In the limit of $n$ applications of a greedification operator $\mathcal{I}$ on a value estimate $q\in\mathcal{Q}$ and a starting policy $\pi_{0}\in\Pi$, the policy $\pi_{n}$ converges to a greedy policy with respect to $q,\,\forall s\in\mathcal{S}$:
$\lim_{n\to\infty}\sum_{a\in\mathcal{A}}q(s,a)\pi_{n}(s)=\max_{a}q(s,a),\quad\text{where}\quad\pi_{n+1}(s)=\mathcal{I}(\pi_{n},q)(s).$ (6)
Intuition.
If a greedification operator cannot converge to an $\operatorname*{arg\,max}$ policy even in the limit for a fixed $Q^{\pi}$, then it is clear that this operator cannot converge to an optimal policy. This is necessitated by the fact that every optimal policy is an $\operatorname*{arg\,max}$ policy. See Supplementary Materials C.1 for a concrete example where such a condition is necessary for convergence of a Policy Iteration algorithm. Since practical operators are not generally designed to distinguish between exact $Q^{\pi}$ and approximated $q\approx Q^{\pi}$, we formulate the definition in terms of $q\in\mathcal{Q}$.
Unfortunately, the necessary greedification condition is not sufficient, even in the case of exact evaluation. This is due to the fact that assuming convergence to a greedy policy in the limit for a fixed $q$ function does not necessarily imply the same when the $q$ function changes between iterations. There exist settings where the ordering of actions $a,a^{\prime},q(s,a)<q(s,a^{\prime})$ can oscillate between iterations, preventing the convergence to greedy policies (See Supplementary Materials C.3 for a concrete example). Below, we identify two additional conditions which are each sufficient for convergence. The first condition resolves this issue by lower-bounding the rate of improvement, which guarantees that the oscillation does not continue infinitely. The second simply augments the necessary greedification condition to require convergence for any sequence of Q functions.
Definition 3 (Bounded Greedification).
We call an operator $\mathcal{I}$ a bounded greedification operator if $\mathcal{I}$ is a greedification operator (Definition 1) and for every $q\in\mathcal{Q}$, $\exists\epsilon>0,$ such that $\forall s\in\mathcal{S}$:
$\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)-\sum_{a\in\mathcal{A}}\pi(a|s)q(s,a)\quad>\quad\epsilon,$
unless $\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)=\max_{a}q(s,a),\,\forall s\in\mathcal{S}$.
Intuition.
The lower bound $\epsilon$ eliminates the possibility of infinitesimal improvements and guarantees convergence to the $\operatorname*{arg\,max}$ policy in finite iterations, preventing the action orderings from oscillating infinitely.
Definition 4 (Limit-Sufficient Greedification).
Let $q_{0},q_{1},\dots\in\mathcal{Q}$ be a sequence of vectors such that $\lim_{n\to\infty}q_{n}=q$ for some $q\in\mathcal{Q}$. Let $\pi_{0},\pi_{1},\dots$ be a sequence of policies where $\pi_{n+1}=\mathcal{I}(\pi_{n},q_{n+1})$ for some operator $\mathcal{I}$. We call an operator $\mathcal{I}$ a Sufficient greedification operator if $\mathcal{I}$ is a greedification operator (Definition 1) and in the limit $n\to\infty$ the improved policy $\pi_{n+1}$ converges to a greedy policy with respect to the limiting value $q,\,\forall s\in\mathcal{S}$:
$\lim_{n\to\infty}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q_{n}(s,a)=\max_{a}q(s,a).$ (7)
Intuition.
Even in the presence of infinitesimal improvement and non-stationary estimates $q_{n}$, the operator is guaranteed to converge to a greedy policy, as long as there exists a limiting value $q$.
Practical operators that are sufficient operators
Bounded greedification is used to establish convergence for MPO (see Appendix A.2, Proposition 3 of (Abdolmaleki et al., 2018)). Similarly, deterministic operators $\mathcal{I}_{det}$ are also bounded greedification operators (see Supplementary Materials C.4 for proof). Bounded greedification operators however cannot contain operators that induce convergence to the greedy policy only in the limit, because the convergence they induce is in finite steps. $\mathcal{I}_{gmz}$ on the other hand induces convergence only in the limit, and in fact is more generally a limit-sufficient greedification operator (see Supplementary Materials C.5 for proof).
The deterministic greedification operator on the other hand does not converge with respect to arbitrary non-stationary sequences $\lim_{n\to\infty}q_{n}$ (see Supplementary Materials C.7), which leads us to conclude that both sets are useful in that they both contain practical operators and neither set contain the other. The greedy operator on the other hand is a member of both sets, demonstrating that the sets are not disjoint either (see Supplementary Materials C.8 for proof).
Equipped with Definitions 3 and 4 we establish our main theoretical result, convergence for both Algorithms 1 and 2 for operators in either set.
Theorem 3 (Convergence of Algorithms 1 and 2).
Generalized Policy Iteration algorithms and their Value Improved extension (Algorithms 1 and 2 respectively) converge for sufficient greedification operators, in finite iterations (for operators defined in Definition 3) or infinite iterations (for operators defined in Definition 4), in finite-horizon MDPs.
Proof sketch: Using induction from terminal states, the proof builds on the immediate convergence of values of terminal states $s_{H}$, convergence of policies at states $s_{H-1}$ and finally on showing that given that $q,\pi$ converge for all states $s_{t+1}$, they also converge for all states $s_{t}$. The evaluation of a greedier policy (line 4 in Algorithm 2) is accepted by the induction that underlies the convergence of Algorithm 1 which allows us to build on the same induction to establish convergence for Algorithm 2. The full proof is provided in the Supplementary Materials. In C.9 for Algorithm 1 with limit sufficient operators and $k=1$, extended to $k\geq 1$ in C.10, to Value-Improved algorithms in C.11, and to bounded operators in C.12.
A corollary of $\mathcal{I}_{gmz}$ being a limit-sufficient greedification operator along with Theorem 3 is the convergence of a process underlying the Gumbel MuZero algorithm. To the best of our knowledge this is the first time convergence has been established for algorithms rooted in this operator.
Corollary 1.
The Generalized Policy Iteration process underlying the Gumbel MuZero algorithm family converges to the optimal policy for finite horizon MDPs, for all $\pi_{0}\in\Pi$ such that $\log\pi(a|s)$ is defined $\forall s\in\mathcal{S},a\in\mathcal{A}$.
In Algorithm 2, the acting policy is decoupled from the evaluated policy. An interesting question is what conditions the operator $\mathcal{I}_{2}$ used to produce the evaluated policy must satisfy in terms of our definitions so far. The following corollary establishes that $\mathcal{I}_{2}$ does not need to be a sufficient or even a necessary greedification operator for convergence to the optimal policy.
Corollary 2.
Algorithm 2 converges to the optimal policy for any non-detriment operator $\mathcal{I}_{2}$ (e.g. operators that satisfy the non-strict inequality of Equation 3), as long as $\mathcal{I}_{1}$ is itself sufficient.
For proof see Supplementary Materials C.11. Motivated that the Generalized Policy Iteration process underlying VIAC algorithms converges, we proceed to evaluate practical VIAC algorithms.
4 Value Improved Actor Critic Algorithms
Value-improvement can be incorporated into existing AC algorithms in one of two ways: (i) Incorporating an additional explicit greedification operator to produce a greedier evaluation policy, and use the greedier policy to bootstrap actions from which to generate value targets. (ii) Incorporating an implicit greedification operator by replacing the value loss with an asymmetric loss (we include pseudo code in Appendix B, Algorithms 3 and 4 respectively and implementation details in Supplementary Materials D).
To verify that value-improvement can be useful in practice, in Figure 1 we evaluate the popular model-free off-policy actor critic algorithms TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018b) against value improved (VI) variants across 12 environments from the DeepMind continuous control benchmark (Tassa et al., 2018). As a value-improvement operator, both VI agents use an implicit improvement operator, replacing the $\mathcal{L}_{2}$ loss of the critic with the asymmetric expectile-loss $\mathcal{L}^{\tau}_{2}$ with $\tau=0.75$ (see Supplementary Materials D.3 for implementation details). Across all environments tested, the VI variation significantly outperforms or matches its respectively baseline, demonstrating that indeed it is possible to benefit from Value Improvement with practical algorithms in standard environments. Because this operator induces implicit improvement, it cannot be applied directly to the acting policy and can only be used for value improvement. In addition, this operators introduces only a negligible increase in compute and implementation cost. In Figure 5 in Appendix A we include additional results for the recent algorithm TD7 (Fujimoto et al., 2023), which show similar gains in this domain. We conclude that value-improvement can be useful in practice.
Figure 1: VI-TD3 and VI-SAC (ours) with expectile loss as a value improvement operator vs. TD3 and SAC respectively, on 12 DeepMind continuous control environments. Mean and two standard errors in the shaded area of evaluation curves across 20 seeds.
Can AC algorithms directly benefit from a greedier evaluated policy (i.e. value-improvement)?
It remains unclear, however, whether the performance gains observed in Figure 1 are a result of increased greedification directly or alternatively due to other beneficial properties of the implicit greedification operator. To evaluate the contribution of greedification as directly as possible, in Figure 2 we compare VI-TD3 where the value improvement operator $\mathcal{I}_{2}=\mathcal{I}_{1}$ is the policy improvement operator used by TD3, the deterministic policy gradient. In order to compare different degrees of greedification, a different number $pg=n$ of repeating gradient steps with respect to the same batch are applied to the evaluated policy, which is then discarded after each use. Increased performance with increased value improvement is demonstrated on the left of Figure 2. In the center we plot the difference between the value bootstrap that uses the greedier policy $\pi^{\prime}$ and the baseline bootstrap, demonstrating that more greedification results in larger (greedier) value targets, as expected. On the right over-estimation bias is evaluated following the method used by Chen et al. (2021): the predicted value is compared to the returns observed in evaluation episodes, averaged across different state-actions and trajectories. The observed over-estimation bias is small (top, hopper-stand) to non-existent (bottom, hopper-hop), suggesting that the value improvement observed in the center is not a result of increased over-estimation bias. This suggestes that TD3 directly benefits from increased greedification of the evaluated policy in this domain.
Figure 2: Mean and one standard error in the shaded area of evaluation curves, across 10 seeds for VI-TD3 with $\mathcal{I}_{2}$ the deterministic policy gradient and increasing number of $n$ gradient steps (pg=n), with baseline (i.e. pg=0) TD3 for reference.
Greedification vs. stability tradeoff
In Figure 3 we evaluate VI-TD3 with increasing values of the greedification-parameter $\tau$. The increased greedification monotonically improves performance up to a point, from which performance monotonically degrades. This suggests that there is a similar tradeoff between greedification and instability in value-improved algorithms and that the tradeoff can directly be optimized by tuning the greediness of the value-improvement operator.
Figure 3: Mean and one standard error across 10 seeds in evaluation for VI-TD3 with expectile loss with different values of the expectile parameter $\tau$. A a monotonic increase in performance in observed up to a point, from which there is a monotonic decrease in performance.
Increased greedification of the acting policy
In Section 1 we motivate that increasing the greediness of the update in a parameterized policy is challenging and that it should not be done by repeating gradient steps on the same batch. We demonstrate this in Figure 4, where it indeed does not improve, and sometimes even degrades, performance. This supports our claim that repeating gradient steps with respect to the same batch is not a good greedification of the acting policy.
Figure 4: Mean and one standard error across 10 seeds in evaluation of VI-TD3 with Policy Gradient as the value improvement operator, vs. TD3 with 20 repeating policy gradient steps in each update, vs. baseline TD3.
Finally, if one is able to spend additional compute on gradient updates, an increased replay ratio is an attractive alternative to value improvement. In Figure 7 in Appendix A we compare VI-TD3 with increasing number of gradient steps to TD3 with increasing replay ratio. In line with similar findings in literature (Chen et al., 2021), replay ratio provides a very strong performance gain for small ratios. As the ratio increases, performance degrades, a result which the literature generally attributes to instability. The VI agent on the other hand does not degrade with increased compute. This suggests a reduced interaction between greedification of the evaluated policy and instability compared to that of the acting policy.
5 Related Work
In model-based RL, employing improvement operators in the form of planning at multiple different steps in the same algorithm is a popular choice (see Moerland et al., 2023). The more common setup employs the same improvement twice: once online to select an action from an improved policy and once during training to improve a parameterized policy (for example, AlphaZero (Silver et al., 2018)). A few algorithms use the same operator a third time to produce an improved evaluation policy on the next state as well, when generating value targets (e.g. MuZero Reanalyze, Schrittwieser et al., 2021). On the other hand, these algorithms can more traditionally be motivated from the perspective that the acting policy, target policy and evaluated policy all coincide as they are all produced by the same operator (MCTS in the case of Reanalyze). From this perspective, these algorithms can be thought of as belonging to the standard AC framework, rather than VIAC. Since the operators used by model based RL algorithms are often very computationally intensive, in practice similar methods propose to drop the improvement to the evaluated policy (see Ye et al., 2021).
In model-free RL, TD3 can be thought of as an example of an agent which acts, improves, and evaluates three different policies: The acting policy is improved using the deterministic policy gradient, during action selection the acting policy is modified with noise in order to induce exploration, and finally the evaluated policy is regularized with a differently-parameterized noise in order to improve learning stability. Although only one policy improvement operator is used, TD3 can be thought of as an algorithm which decouples the acting policy from the evaluated policy. GreedyAC (Neumann et al., 2023) shares similarities with the VIAC framework in that it explicitly maintains two different policies, one more and one less greedy. Both policies are used during the same policy-improvement step however using a conditional-cross-entropy method, and the evaluated policy remains the acting policy, as in standard AC methods.
6 Conclusions
In order to better control the tradeoff between greedification and stability in AC algorithms we propose to decouple the evaluated policy from the acting policy and apply a policy improvement step additionally to the evaluated policy. Since this improvement is retained only in the value function we refer to this approach as Value-Improved AC (VIAC). We identify sets of operators for which a Dynamic Programming process underlying this approach, Value-Improved Generalized Policy Iteration, converges. We demonstrate that policy improvement itself is not a sufficient condition for convergence of Dynamic Programming algorithms with stochastic policies. We identify necessary and sufficient conditions for convergence of such algorithms, and prove that Generalized Policy Iteration algorithms converge to the optimal policy for such sufficient greedification operators in the finite horizon domain. We prove that the greedification operator used by the Gumbel MuZero algorithm is an example of a sufficient greedification operator. As a corollary, this establishes that a Generalized Policy Iteration process underlying the Gumbel MuZero algorithm family similarly converges. Empirically, VI-TD3 and VI-SAC significantly improve upon or match the performance of their respective baselines in all DeepMind control environments tested with negligible increase in compute and implementation costs. Our experiments further suggest that the controllable greediness of the evaluated policy, which becomes possible with value-improvement, directly controls a tradeoff between greedification and stability. We hope that our work will act as motivation to design future Actor Critic algorithms with multiple improvement operators in mind, as well as extend existing algorithms with a value-improvement step.
Appendix A Additional Results
A.1 Value Improved TD7
Figure 5: Mean and two standard errors across 10 seeds of VI-TD7 with expectile loss vs. TD7 on the same tasks as Figure 1. Similar performance gains are observed for VI-TD7 in this domain.
A.2 Greedification with sampling based policies
Figure 6: Mean and one standard error in the shaded area of evaluation curves, across 10 seeds for VI-TD3 with gradient, sampling and implicit value improvement operators $\mathcal{I}_{2}$. Sampling-based operators allow for explicit greedification without a significant increase in sequential compute, compared to gradient based greedification operators. Similar performance gains are demonstrated by the sampling based operator compared to the gradient based, while implicit remains the best performer.
A.3 Increased value improvement vs. increased replay ratio
Figure 7: Mean and one standard errors across 10 seeds in evaluation of VI-TD3 with policy-gradient based value improvement vs. td3 with increased replay ratio. Number of gradient steps are equated across rr / VI agent pairs as a pseudo metric for compute. The performance of TD3 generally degrades with increased replay ratio, in line with the results of Chen et al. (2021). In contrast, the performance of VI-TD3 increases with compute, without access to additional mechanisms to address instability.
Appendix B Explicit and implicit Value Improved Actor Critic algorithms
In Algorithms 3 and 4 on Page 3, modifications to baseline off-policy Actor Critic are marked in blue.
Algorithm 3 Explicit Off-policy Value-Improved Actor Critic
1:Initialize policy network $\pi_{\theta}$, $Q$ network $q_{\phi}$, Greedification Operators $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, replay buffer $\mathcal{B}$
2:for each episode do
3:     for each environment interaction $t$ do
4:         Act $a_{t}\sim\pi_{\theta}(s_{t})$
5:         Observe $s_{t+1},r_{t}$
6:         Add the transition $(s_{t},a_{t},r_{t},s_{t+1})$ to the buffer $\mathcal{B}$
7:         Sample a batch $b$ from $\mathcal{B}$ of transitions of the form $(s_{t},a_{t},r_{t},s_{t+1})$
8:         Update the policy $\pi_{\theta}(s_{t})\leftarrow\mathcal{I}_{1}(\pi_{\theta},q_{\phi})(s_{t}),\forall s_{t}\in b$
9:         Further improve the policy $\pi^{\prime}(s_{t+1})\leftarrow\mathcal{I}_{2}(\pi_{\theta},q_{\phi})(s_{t+1}),\forall s_{t+1}\in b$
10:         Sample an action from the improved policy $a\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pi^{\prime}}(s_{t+1}),\forall s_{t+1}\in b$
11:         Compute the value targets $y(s_{t},a_{t})\leftarrow r_{t}+\gamma q_{\phi}(s_{t+1},a),\forall(s_{t},a_{t},r_{t},s_{t+1})\in b$
12:         Update $q_{\phi}$ with gradient descent and MSE loss using targets $y$
Algorithm 4 Implicit Off-policy Value-Improved Actor Critic
1:Initialize policy network $\pi_{\theta}$, $Q$ network $q_{\phi}$, Greedification Operator $\mathcal{I}_{1}$, implicit greedification parameter $\tau$ and replay buffer $\mathcal{B}$
2:for each episode do
3:     for each environment interaction $t$ do
4:         Act $a_{t}\sim\pi_{\theta}(s_{t})$
5:         Observe $s_{t+1},r_{t}$
6:         Add the transition $(s_{t},a_{t},r_{t},s_{t+1})$ to the buffer $\mathcal{B}$
7:         Sample a batch $b$ from $\mathcal{B}$ of transitions of the form $(s_{t},a_{t},r_{t},s_{t+1})$
8:         Update the policy $\pi_{\theta}(s_{t})\leftarrow\mathcal{I}_{1}(\pi_{\theta},q_{\phi})(s_{t}),\forall s_{t}\in b$
9:         Sample an action from the policy $a\sim\pi(s_{t+1}),\forall s_{t+1}\in b$
10:         Compute the value targets $y(s_{t},a_{t})\leftarrow r_{t}+\gamma q_{\phi}(s_{t+1},a),\forall(s_{t},a_{t},r_{t},s_{t+1})\in b$
11:         Update $q_{\phi}$ with gradient descent and $\mathcal{L}^{\tau}_{2}$ loss using targets $y$, see Supplement D.3
Acknowledgments
We acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc) as well as the DAIC cluster. This work was partially supported by the EU Horizon 2020 programme under grant number 964505 (Epistemic AI), and partially funded by the Dutch Research Council (NWO) project Reliable Out-of-Distribution Generalization in Deep Reinforcement Learning with project number OCENW.M.21.234.
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix C Proofs
C.1 Theorem 2: policy improvement is not enough
Theorem 2 states: Policy improvement is not a sufficient condition for the convergence of Policy Iteration algorithms (Algorithm 1 with exact evaluation) to the optimal policy for all starting policies $\pi_{0}\in\Pi$ in all finite-state MDPs.
Proof sketch: We will construct a simple MDP where the $Q^{\pi}$ values remain the same for all policies $\pi$, and show that even in this simple example, it is possible for a Policy Iteration algorithm to converge to non-optimal policies, with policy improvement operators that allow for stochastic policies. In addition, while adjacent to the narrative of this paper, we demonstrate that the same problem persists with deterministic policies in continuous action spaces in Appendix C.1.1.
Proof.
Consider a very simple deterministic MDP with starting state $s_{0}$ and two actions $a_{1},a_{2}$ that lead respectively to two terminal states $s_{1},s_{2}$. The reward function $R(s_{0},a_{1})=1$ and $R(s_{0},a_{2})=2$ and transition function $P(s_{1}|s_{0},a_{1})=1,P(s_{2}|s_{0},a_{2})=1$ and zero otherwise. We have $Q^{\pi}(s_{0},a_{1})=1$ and $Q^{\pi}(s_{0},a_{2})=2$ for all policies $\pi\in\Pi$. The optimal policy is therefor $\pi^{*}(s_{0})=a_{2}$, that is $\pi^{*}(a_{1}|s_{0})=0$ and $\pi^{*}(a_{2}|s_{0})=1$.
Consider the very simple policy improvement operator $\mathcal{I}_{inadequate}$, defined as follows: When $\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}(s,a)<\mathrm{softmax}(Q^{\pi})(a|s)Q^{\pi}(s,a)$, $\mathcal{I}_{inadequate}(\pi)=\alpha\pi+(1-\alpha)\mathrm{softmax}(Q^{\pi})$. When $\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}(s,a)\geq\mathrm{softmax}(Q^{\pi})(a|s)Q^{\pi}(s,a)$, $\mathcal{I}_{inadequate}(\pi)=\operatorname*{arg\,max}_{a}Q^{\pi}$.
In natural language: when the policy is less greedy than the softmax policy, the operator produces a mix between the current policy and the softmax policy. This is always greedier than the current policy, and thus will act as a greedification operator for policies such policies. When the policy is as greedy or greedier than the softmax, the operator produces directly a greedy policy.
The policy improvement theorem proves that this operator is a Policy Improvement operator, because for every policy $\pi$ it greedifies the policy with respect to $Q^{\pi}$.
We now apply this operator in the Policy Iteration scheme to the simple example MDP specified above, with a starting uniform policy $\pi(a_{1}|s_{0})=\pi(a_{2}|s_{0})$. Since this operator produces a mixture of the current and softmax policy, in the limit, it will converge to the softmax policy, i.e. to a sub-optimal policy, even though it is a policy improvement operator. ∎
In this example, since $Q^{\pi}$ does not change across different iterations $n$ of a Policy Iteration algorithm, we can identify the following: For convergence to the optimal policy, it is necessary that a greedification operator $\pi_{n+1}=\mathcal{I}(\pi,q)$ will converge to a greedy policy, with respect to the same stationary $q=Q^{\pi}$:
$\lim_{n\to\infty}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)=\max_{a}q(s,a),\,\forall s\in\mathcal{S}.$
C.1.1 Policy improvement in continuous action spaces
[…]
Let us again choose $\pi_{0}$ the uniform policy across actions. At the first step, $\mathcal{I}$ can produce just-above the middle action $a_{1}>0.5$. At each step, $\mathcal{I}$ can produce a new action $\mathcal{I}(\pi_{1},Q^{*})(s_{0})=\pi_{2}(s_{0})=a_{2}>a_{1}$. However, since the space $(0,1)$ is the continuum and non-countable, there are more actions to select that any iterative process will ever have to go through. Therefor, even in the limit $n\to\inf$, the operator will never have to choose $\mathcal{I}(\pi_{n},Q^{*})=1$.
C.2 Policy Improvement operators that are not Greedification operators
Lemma 1.
There exist operators $\mathcal{I}:\Pi\times\mathcal{Q}\to\Pi$ that are Policy Improvement operators, and therefore fulfill Equation 2, but are not Greedification operators according to Definition 1.
Proof sketch: We will prove that a random-search operator that mutates the policy $\pi$ randomly into a new policy $\pi^{\prime}$, evaluates $\pi^{\prime}$ and keeps it if $V^{\pi}>V^{\pi^{\prime}}$, is not a greedification operator, even though it is a policy improvement operator by definition. We do this by constructing an MDP and choosing a specific initial policy $\pi_{0}$. Greedification with respect to the initial policy, at state $s_{0}$, results in $\pi_{1}(s_{0})=a_{1}$. However, the optimal policy in this state actually chooses action $a_{0}$, because the optimal policy can take better actions in the future than policy $\pi_{1}$. Such an example proves that it is possible to construct policy improvement while violating greedification, demonstrating that the condition only goes one way: every greedification operator is policy improvement operator, not vice versa.
Proof.
Consider the following finite-horizon MDP: State space $\mathcal{S}=\{s_{1},\dots,s_{10}\}$. Action space $\mathcal{A}={a_{1},a_{2},a_{3}}$. States $\{s_{5},\dots,s_{10}\}$ are terminal states. Transition function: $f(s_{1},a_{1})=s_{2}$, $f(s_{1},a_{2})=s_{3}$, $f(s_{1},a_{3})=s_{4}$, $f(s_{2},a_{1})=s_{5}$, $f(s_{2},a_{2})=s_{6}$, $f(s_{3},a_{1})=s_{7}$, $f(s_{3},a_{2})=s_{8}$, $f(s_{4},a_{1})=s_{9}$, $f(s_{4},a_{2})=s_{10}$.
Rewards: $R(s_{2},s_{5})=2$, $R(s_{2},s_{6})=-1$, $R(s_{3},s_{7})=1$, $R(s_{3},s_{8})=0$, $R(s_{4},s_{9})=3$, $R(s_{4},s_{10})=-2$.
Actions that are not specified lead directly to a terminal state with zero reward.
Let us begin by identifying the optimal policy in this MDP, in states $s_{1}$ and $s_{4}$: $\pi^{*}(s_{1})=a_{3}$ and $\pi^{*}(s_{4})=a_{1}$, with a value of $3$ without a discount factor.
Let us construct a starting policy $\pi_{0}$:
$\pi_{0}(s_{1})=a_{1},\,\pi_{0}(s_{2})=a_{2},\,\pi_{0}(s_{3})=a_{2},\pi_{0}(s_{4})=a_{2}$. The other states are terminal and there are no actions to take, and therefor no policy.
Consider the following Policy Improvement operator $\mathcal{I}_{E}:\Pi\times\mathcal{Q}\to\Pi$, which this example will demonstrate is not a greedification operator. $\mathcal{I}_{E}$ takes a policy $\pi$, and mutates it with a random process to $\pi^{\prime}$. $\mathcal{I}_{E}$ proceeds to conduct exact evaluation of $\pi^{\prime}$, to find $V^{\pi^{\prime}}$. If $V^{\pi^{\prime}}(s)\geq V^{\pi}(s)$ on all states, and $V^{\pi^{\prime}}(s)>V^{\pi}(s)$ in at least one state, $\mathcal{I}_{E}$ outputs $\pi^{\prime}$. Otherwise, the process repeats. This process guarentees policy improvement. However, this process may directly produce the optimal policy in this MDP, which in states $s_{1},s_{3}$ is $\pi^{*}(s_{1})=a_{3}$ and $\pi^{*}(s_{4})=a_{1}$.
Note however, that the optimal policy is not a greedier policy with respect to the value of $\pi_{0}$. For $\pi_{0}$, we have: $Q^{\pi_{0}}(s_{1},a_{1})=-1,Q^{\pi_{0}}(s_{1},a_{2})=0,Q^{\pi_{0}}(s_{1},a_{3})=-2$. A greedier policy with respect to these values cannot deterministically choose action $a_{3}$, which is the action chosen by the optimal policy in this state.
Therefor, this example demonstrates that it is possible for a policy to be improved (higher value in at least one state, and greater or equal on all states), without being greedier with respect to some original policy’s value. In turn, this demonstrates that there exist Policy Improvement operators that are not Greedification operators. ∎
C.3 Necessary greedification operators may not be sufficient
Lemma 2.
Greedification operators (Definition 1) which have the necessary greedification property (Definition 2) may not be sufficient for Policy Iteration algorithms to converge to the optimal policy.
Proof sketch: First, we demonstrate the problem: certain operators with the necessary property, such as the deterministic greedification operators, may not converge with respect to non-stationary $q_{n}$. Second, we will show that this can happen in practice, even in (exact) Policy Iteration, by constructing an operator that performs deterministic greedification in some states, and greedification that converges only in the limit in other states, and show that the problem can persist in practical MDPs.
Proof.
Let $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and a sequence $q_{n}(a_{1})=(-1)^{n}/2^{n}+q(a_{1})$, $q_{n}(a_{2})=(-1)^{n+1}/2^{n}+q(a_{2})$, and $q_{n}(a_{3})=q(a_{3})$, with a limiting value $q=[1,1,2]$. We omit the dependency of $q$ on a state as it is unnecessary in this example. In this case, the optimal policy with respect to any $q_{n}$ is $\pi=a_{3}$.
Take the least-greedifynig deterministic greedification operator $\mathcal{I}_{min\_det}(q,\pi)=\min_{q(a)>q(\pi),a\neq\pi}q(a)$. This operator produces the worst action, with respect to $q$, that is better than the current action selected by the policy, and as such, is a greedification operator by definition, with respect to deterministic policies. Since there are finitely many deterministic policies on a bounded action space $|\mathcal{A}|<\infty$, this operator will converge to $\lim_{n\to\infty}\pi_{n}=\operatorname*{arg\,max}_{a}q(a)$ with respect to a stationary q.
Take $\pi_{0}=a_{2}$. Using the operator $\mathcal{I}_{min\_det}$ we have $\pi_{n}=\mathcal{I}_{min\_det}(q_{n},\pi_{n-1})$. When $n$ is odd, $\pi_{n}=a_{1}$, and when $n$ is even, $\pi_{n}=a_{2}$, without ever converging to the optimal policy $\pi=a_{3}$.
Next we will construct an example MDP and improvement operators in which this situation can happen in practice. Consider a finite-state, finite horizon MDP with states $s_{1},\dots,s_{n}$. We are interested in the behavior at state $s_{0}$ specifically, which similarly has actions $a_{1},a_{2},a_{3}$, with rewards $R(s_{1},a_{3})=3,R(s_{1},a_{1})=R(s_{1},a_{2})=0$. The transition $f(s_{1},a_{3})=s_{0}$ is terminal and $f(s_{1},a_{1})=s_{2}$, $f(s_{1},a_{2})=s_{3}$.
Consider the following improvement operator: On state $s_{1}$, this operator is $\mathcal{I}_{min_{d}et}$. However, on all other states, this is a necessary greedification operator, which converges only in the limit, and in a non-constant rate. It is possible to construct the rest of the MDP and starting policies $\pi_{0}$ such that the sequence alternates $1>Q^{\pi_{n}}(s_{1},a_{1})>Q^{\pi_{n}}(s_{1},a_{2})$ when $n$ is odd, and $1>Q^{\pi_{n}}(s_{1},a_{2})>Q^{\pi_{n}}(s_{1},a_{1})$ when $n$ is even, while both are smaller than $Q^{\pi_{n}}(s_{1},a_{3})=Q^{*}(s_{1},a_{3})=3$. This is possible because the policies $\pi_{n}(s_{2}),\pi_{n}(s_{3})$ can be soft, and it is possible to construct an MDP which produces arbitrary values bounded between $0,1$ by setting $R(s_{2},a_{1})=1,R(s_{2},a_{2})=0,R(s_{2},a_{3})=0$ and $R(s_{3},a_{1})=1,R(s_{3},a_{2})=0,R(s_{3},a_{3})=0$. In such MDP, $\lim_{n\to\infty}\pi_{n}(s_{1})$ will never converge to $a_{3}$, the optimal policy in this state. ∎
C.4 Deterministic greedification operators are bounded-greedification operators
Lemma 3.
Deterministic greedification operators $\mathcal{I}_{det}$, i.e. greedification operators (Definition 1) that produce deterministic policies are bounded greedification operators 3.
Proof.
Take $\epsilon=\min_{s\in\mathcal{S},a,a^{\prime}\in\mathcal{A},q(s,a)\neq q(s,a^{\prime})}|q(s,a)-q(s,a^{\prime})|$, that is, the minimum difference across two actions that do not have the same value (i.e. the minimum greater than zero difference). If there is no greater than zero difference, then all actions are optimal and every policy is already optimal. Otherwise, the greedification imposed by choosing at least one better action in at least one state has to be greater than the minimum difference between two actions. ∎
C.5 The operator $\mathcal{I}_{gmz}$ is a Limit-Sufficient Greedification operator
The operator proposed by Danihelka et al. (2022) is defined as follows:
$\mathcal{I}_{gmz}(\pi,q)(a|s)=\mathrm{softmax}(\sigma(q(s,a))+\log\pi(a|s))={\textstyle\frac{\exp(\log\pi(a|s)+\sigma(q(s,a)))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\log\pi(a^{\prime}|s)+\sigma(q(s,a^{\prime})))}}$ (8)
Lemma 4 ($\mathcal{I}_{gmz}$ with a stationary $\sigma$ is a Limit-Sufficient Greedification Operator).
For any starting policy $\pi_{0}\in\Pi$ such that $\log\pi_{0}(a|s)$ is defined and sequences $q_{1},\dots,q_{n}$ such that $\lim_{n\to\infty}q_{n}=q\in\mathcal{Q}$, iterative applications $\pi_{n+1}=\mathcal{I}_{gmz}(\pi_{n},q_{n})$ converge to a greedy policy with respect to the limiting value $q$.
That is,
$\lim_{n\to\infty}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q_{n}(s,a)=\max_{b}q(s,b),\quad\forall s\in\mathcal{S}.$
Proof sketch: We will prove that $n$ repeated applications of the $\mathcal{I}_{gmz}$ operator converge to a softmax policy of the form
$\pi_{n}(a|s)\propto\exp(\log\pi_{0}(a|s)+n\sigma(q_{n}(s,a))),$
which itself converges to an $\operatorname*{arg\,max}$ policy as $\lim_{n\to\infty}$. For simplicity, we will first prove for a stationary $q$, and then repeat the same steps for a non-stationary $q_{n},\lim_{n\to\infty}=q$ for some limiting value $q$.
Proof.
We will show that the Gumbel MuZero operator $\mathcal{I}_{gmz}$ with $\sigma$ a monotonically increasing transformation, is a Limit-Sufficient Greedification operator.
Danihelka et al. (2022) have shown that this operator is a Greedification operator (Section 4 and Appendix C of (Danihelka et al., 2022)). It remains for us to demonstrate that the sequence $\pi_{n}$ converges for $\mathcal{I}_{gmz}$, such that
$\lim_{n\to\infty}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)=\max_{b}q(s,a),$
for any $\pi_{0}$ and $\forall s\in\mathcal{S}$.
Step 1: Convergence with stationary $q$
For a stationary $q$, at any iteration $n$, the policy $\pi_{n}$ can be formulated as:
$\pi_{n}(a)=\frac{1}{z_{n}}\exp(\sigma(q(s,a))+\log\pi_{n-1}(a|s)),\quad z_{n}=\sum_{a^{\prime}\in\mathcal{A}}\exp(\sigma(q(s,a^{\prime}))+\log\pi_{n-1}(a^{\prime}|s))$ (9)
Where $z_{n}$ is the normalizer of the $\mathrm{softmax}$ operator. We can expand $\pi_{n}$ backwards as follows:
$\pi_{n}(a)$ $=\frac{1}{z_{n}}\exp(\sigma(q(s,a))+\log\pi_{n-1}(a|s))$ (10)
$=\frac{1}{z_{n}}\exp\big{(}\sigma(q(s,a))+\log\frac{\sigma(q(s,a))+\pi_{n-1}(a|s)}{z_{n-1}}\big{)}$ (11)
$=\frac{1}{z_{n}}\exp\big{(}\sigma(q(s,a))+\sigma(q(s,a))+\log\pi_{n-2}(a|s)-\log z_{n-1}\big{)}$ (12)
$=\frac{1}{z_{n}z_{n-1}}\exp\big{(}2\sigma(q(s,a))+\log\pi_{n-2}(a|s)\big{)}$ (13)
=$\dots$= … (14)
$=(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\big{(}n\sigma(q(s,a))+\log\pi_{0}(a|s)\big{)}$ (15)
As $\pi_{n}$ is a softmax policy, i.e. $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)=1$, the product $\Pi_{i=1}^{n}\frac{1}{z_{i}}$ must act as a normalizer:
$\Pi_{i=1}^{n}\frac{1}{z_{i}}=\sum_{a\in\mathcal{A}}\exp\big{(}n\sigma(q(s,a))+\log\pi_{0}(a|s)\big{)}$ (16)
We can now directly take the limit $\lim_{n\to\infty}\pi_{n}$:
$\lim_{n\to\infty}\pi_{n}(a)=\lim_{n\to\infty}(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\big{(}n\sigma(q(s,a))+\log\pi_{0}(a|s)\big{)}$ (17)
It is well established that as the temperature $1/n$ goes to zero, the $\mathrm{softmax}$ converges to an $\operatorname*{arg\,max}$ (Collier et al., 2020). With non-stationary $q_{n}$ we get a slightly more involved sequence, and the formulated proof that the $\mathrm{softmax}$ converges to an $\operatorname*{arg\,max}$ will serve us to demonstrate convergence with $q_{n}$. We include the proof that the $\mathrm{softmax}$ converges to an $\operatorname*{arg\,max}$ below in step 1.5.
Step 1.5: Convergence of the $\mathrm{softmax}$ to an $\operatorname*{arg\,max}$
Define $\sigma_{max}=\max_{a}\sigma(q(s,a))$. Let us now multiply by $\frac{\exp(-n\sigma_{max})}{\exp(-n\sigma_{max})}$. We have:
$\pi_{n}(a|s)$ $=(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\big{(}n\sigma(q(s,a)+\log\pi_{0}(a|s))\big{)}\frac{\exp(-n\sigma_{max})}{\exp(-n\sigma_{max})}$ (18)
$=\frac{\pi_{0}(a|s)}{\exp(-n\sigma_{max})}(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\Big{(}n\big{(}\sigma(q(s,a))-\sigma_{max}\big{)}\Big{)}$ (19)
We now note that $\sigma(q(s,a))-\sigma_{max}<0$ if $\sigma(q(s,a))\neq\max_{a}\sigma(q(s,a))=\sigma_{max}$ and otherwise $\sigma(q(s,a))-\sigma_{max}=0$ if $\sigma(q(s,a))=\max_{a}\sigma(q(s,a))=\sigma_{max}$. In that case, $\exp\big{(}n(\sigma(q(s,a))-\sigma_{max})\big{)}=\exp\big{(}0\big{)}=1$. We substitute that into the limit:
$\lim_{n\to\infty}\pi_{n}(a|s)=\begin{cases}\lim_{n\to\infty}\frac{\pi_{0}(a|s)}{\exp(-n\sigma_{max})}(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\Big{(}n\big{(}\sigma(q(s,a))-\sigma_{max}\big{)}\Big{)}=0,&\text{if }\sigma(q(s,a))\neq\sigma_{max}\\ \lim_{n\to\infty}\frac{\pi_{0}(a|s)}{\exp(-n\sigma_{max})}(\Pi_{i=1}^{n}\frac{1}{z_{i}})1,&\text{if }\sigma(q(s,a))=\sigma_{max}\end{cases}$ (20)
Note that:
  1. 1.
    The numerator where $\sigma(q(s,a))=\sigma_{max}$ converges to $e^{0}=1$
  2. 2.
    The numerator where $\sigma(q(s,a))\neq\sigma_{max}$ converges to $\lim_{n\to\infty}e^{-\delta n}=0$, $\delta>0$.
  3. 3.
    The denominator always normalizes the policy such that $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)=1,\forall s\in\mathcal{S}$, due to the definition of the softmax.
As a result, we have:
$\lim_{n\to\infty}\pi_{n}(a|s)=\begin{cases}\frac{0}{z},&\text{if }\sigma(q(s,a))\neq\sigma_{max}\\ \frac{1}{z},&\text{if }\sigma(q(s,a))=\sigma_{max}\end{cases}$ (21)
For some normalization constant $z$. I.e. the policy $\lim_{n\to\infty}\pi_{n}$ is an $\operatorname*{arg\,max}$ policy with respect to $q$, that is, the policy has probability mass only over actions that maximize $\sigma(q)$.
Step 2: Convergence with non-stationary $q_{n}$
We will now extend the proof to a non-stationary $q_{n}$ that is assumed to have a limiting value, $\lim_{n\to\infty}q_{n}=q$, in line with definition of Sufficient Greedification.
First, we have:
$\pi_{n}(a)$ $=\frac{1}{z_{n}}\exp(\sigma(q_{n}(s,a))+\log\pi_{n-1}(a|s))$ (22)
$=\frac{1}{z_{n}}\exp\big{(}\sigma(q_{n}(s,a))+\log\frac{\sigma(q_{n-1}(s,a))+\pi_{n-1}(a|s)}{z_{n-1}}\big{)}$ (23)
$=\frac{1}{z_{n}z_{n-1}}\exp\big{(}\sigma(q_{n}(s,a))+\sigma(q_{n-1}(s,a))+\log\pi_{n-2}(a|s)\big{)}$ (24)
$=(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\Big{(}\big{(}\sum_{i=1}^{n}\sigma(q_{i}(s,a))\big{)}+\log\pi_{0}(a|s)\Big{)}$ (25)
Based on the same expansion of the sequence as above. Multiplying by $\frac{-n\sigma_{max}}{-n\sigma_{max}}$ and formulating the limit in a similar manner to above, we then have:
$\lim_{n\to\infty}\pi_{n}(a|s)=\lim_{n\to\infty}\frac{\pi_{0}(a|s)}{-n\sigma_{max}}\exp(\Pi_{i=1}^{n}\frac{1}{z_{i}})\big{(}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})\big{)}$ (26)
Let us look at the term $\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})$. First, where $\sigma(q(s,a))\neq\sigma_{max}$, we have
$\lim_{n\to\infty}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})=\lim_{n\to\infty}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma(q(s,a))-(\sigma_{max}-\sigma(q(s,a)))$ (27)
$=\lim_{n\to\infty}-n(\sigma_{max}-\sigma(q(s,a)))+\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma(q(s,a))$ (28)
As the term $\sigma(q_{n}(s,a))-\sigma(q(s,a)$ goes to zero due to the definition of $q_{n}$, the term $\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma(q(s,a))$ goes to a constant, and the term $-n(\sigma_{max}-\sigma(q(s,a)))$ goes to $\infty$- ∞ due to the definition of $\sigma_{max}$. Therefor, the limit:
$\lim_{n\to\infty}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})=-\infty\quad\Rightarrow\quad\lim_{n\to\infty}\exp\big{(}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})\big{)}=0$ (29)
Let us look at the second case, where $\sigma(q(s,a))=\sigma_{max}$, the sequence converges:
$\lim_{n\to\infty}\sum_{i=1}^{n}(\sigma(q_{i}(s,a))-\sigma_{max})=\alpha(s,a)$ (30)
For some constant $\alpha(s,a)$, as $\lim_{n\to\infty}\sigma(q_{n}(s,a))=\sigma_{max}$. Thus, we have again:
$\lim_{n\to\infty}\pi_{n}(a|s)=\begin{cases}\frac{0}{z},&\text{if }\sigma(q(s,a))\neq\sigma_{max}\\ \frac{\alpha(s,a)}{z},&\text{if }\sigma(q(s,a))=\sigma_{max}\end{cases}$ (31)
Demonstrating that $\pi_{n}$ converges to an $\operatorname*{arg\,max}$ policy with respect to $\sigma(q)$. Since $\sigma$ is monotonically increasing, $q(s,a)$ and $\sigma(q(s,a))$ are maximized for the same action $a$, thus $\pi_{n}$ is also an $\operatorname*{arg\,max}$ policy with respect to $q$. Therefor, $\mathcal{I}_{gmz}$ is a Limit-Sufficient Greedification operator. ∎
C.6 $\mathcal{I}_{gmz}$ can be formulated as an insufficient-greedification operator
Lemma 5.
The $\mathcal{I}_{gmz}$ greedification operator with a non-stationary $\sigma$ transformation can be formulated as an insufficient greedification operator.
Proof sketch: We construct a variation of the $\mathcal{I}_{gmz}$ operator with an increasing transformation $\sigma_{n}$, which is different at each iteration. Because the transformation is not constant, it converges to some $\mathrm{softmax}$ policy rather than an $\operatorname*{arg\,max}$ policy.
Proof.
The function $\sigma$ used by $\mathcal{I}_{gmz}$ is only required to be an increasing transformation (see Danihelka et al. (2022), Section 3.3). That is if $q(s,a)>q(s,a^{\prime})$ then we must have that $\sigma(q(s,a))>\sigma(q(s,a^{\prime}))$. In practice, the function proposed by Danihelka et al. (2022) is of the form $\sigma(q(s,a))=\beta(N)q(s,a)$, where $\beta$ is a function of the planning budget $N$ of the MCTS algorithm.
A practitioner might be interested in running the algorithm with a decreasing planning budget over iterations (perhaps the value estimates become increasingly more accurate, and therefor there is less reason to dedicate much compute into planning with MCTS). In that case, we can formulate $\sigma_{n}(q_{n}(s,a))=\frac{\alpha}{\beta^{n}}q_{n}(s,a)$. This transformation is always increasing in $q(s,a)$, adhering to the requirements from $\sigma$. Nonetheless, the sequence $\pi_{n}$ will not converge to an argmax policy for this choice of $\sigma$:
$\lim_{n\to\infty}\pi_{n}$ $=\lim_{n\to\infty}(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\big{(}\sum_{i\leq n}\big{[}\sigma_{n}(q_{n}(s,a))\big{]}+\log\pi_{0}(a|s)\big{)}$ (32)
$=\lim_{n\to\infty}(\Pi_{i=1}^{n}\frac{1}{z_{i}})\exp\big{(}\frac{\alpha}{\beta^{n}}\sum_{i\leq n}\big{[}q_{n}(s,a)\big{]}+\log\pi_{0}(a|s)\big{)}$ (33)
Which will converge to some softmax policy as the following limit converges to a constant: $\lim_{n\to\infty}\frac{\alpha}{\beta^{n}}\sum_{i\leq n}\big{[}q_{n}(s,a)\big{]}=c(s,a),$ and thus the policy remains a softmax policy $\pi_{n}(a|s)=\mathrm{softmax}(c(s,a)+\log\pi_{0}(a|s))$. ∎
C.7 Bounded Greedification operators $\not\subset$ Limit-Sufficient Greedification operators
Lemma 6.
The set of all bounded greedification operators (Definition 3) is not a subset of the set of all limit-sufficient greedification operators (Definition 4). That is, there exists a bounded greedification operator which is not a limit-sufficient greedification operator.
Proof sketch: Convergence with respect to arbitrary sequences $\lim_{n\to\infty}q_{n}=q$ is a strong property, and it is possible to come up with sequences for which specific Bounded Greedification operator do not result in convergence. By constructing such a sequence and choosing such an operator, we will show that there are Bounded-Greedification operators which are not Limit-Sufficient Greedification operators, demonstrating that Bounded Greedification operators $\not\subset$ Limit-Sufficient Greedification operators.
Proof.
Let $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and a sequence $q_{n}(a_{1})=(-1)^{n}/2^{n}+q(a_{1})$, $q_{n}(a_{2})=(-1)^{n+1}/2^{n}+q(a_{2})$, and $q_{n}(a_{3})=q(a_{3})$, with a limiting value $q=[1,1,2]$. A Limit-Sufficient Greedification operator operating on this sequence $\pi_{n+1}=\mathcal{I}_{s}(\pi_{n},q_{n})$ will converge to a greedy policy $\pi_{n}=a_{3}$.
On the other hand, the minimal deterministic Greedification operator $\mathcal{I}_{det}(\pi,q)(s)=\operatorname*{arg\,min}_{a}q(s,a)>\sum_{a^{\prime}\in\mathcal{A}}\pi(a^{\prime}|s)q(s,a^{\prime})$, that is, the deterministic Greedification operator which chooses the least-greedifying action at each step will not converge to the optimal policy on this sequence. At each iteration, $\mathcal{I}_{det}(q,\pi_{n})=a_{1,2}$ (as in, $a_{1}$ or $a_{2}$), because $q_{n}$ alternates $q_{n}(a_{1})>q_{n}(a_{2})$ for even $n$, and $q_{n}(a_{1})<q_{n}(a_{2})$ for odd $n$. Since this operator is a bounded greedification operator (see Appendix C.4), this demonstrates that bounded greedification operators $\not\subset$ limit-sufficient greedification operators. ∎
C.8 The greedy operator is both a limit-sufficient as well as a bounded greedification operator
Lemma 7.
The greedy operator $\mathcal{I}_{\operatorname*{arg\,max}}$ is both a bounded greedification operator (Definition 3) as well as a limit-sufficient greedification operator (Definition 4).
Proof.
The greedy operator is a greedification operator by definition. We will show that it can have both the bounded greedification property as well as the limit sufficient greedification property.
Step 1): We will show that the greedy operator is a bounded greedification operator (Definition 3).
The greedy operator produces the maximum greedification in any state by definition. Therefor:
$\sum_{a\in A}\mathcal{I}_{argmax}(\pi,q)(a|s)q(s,a)\geq\mathcal{I}_{det}(\pi,q)(a|s)q(s,a),$
where $I_{det}$ is the deterministic greedification operator, $\forall s\in\mathcal{S},a\in\mathcal{A}$. Since the deterministic greedification operator is itself bounded by an $\epsilon$ (see Appendix C.4), we have $|\sum_{a\in A}\mathcal{I}_{argmax}(\pi,q)(a|s)q(s,a)-\sum_{a\in A}\pi(a|s)q(s,a)|>\epsilon$.
Step 2): We will show that the greedy operator is a limit-sufficient greedification operator (Definition 4).
We will prove that the sequence $(\pi_{n},q_{n})$ defined for $\mathcal{I}_{\operatorname*{arg\,max}}$ as above converges, such that $\lim_{n\to\infty}|\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q_{n}(s,a)-\max_{b}q(s,b)|=0$, for any $\pi_{0}$. That is, the policy converges to an $\operatorname*{arg\,max}$ policy with respect to the limiting value $q$.
For any $q_{n}$ in the sequence, we have by definition of the operator $\sum_{a\in\mathcal{A}}\mathcal{I}_{\operatorname*{arg\,max}}(q_{n},\pi_{n-1})(a|s)q_{n}(s,a)=\max_{a}q_{n}k(s,a)$. We can substitute that into the limit:
$\lim_{n\to\infty}|\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q_{n}(s,a)-\max_{b}q(s,b)|$ (34)
$=\lim_{n\to\infty}|\max_{a}q_{n}(s,a)-\max_{b}q(s,b)|$ (35)
$\leq\lim_{n\to\infty}\max_{a}|q_{n}(s,a)-q(s,a)|$ (36)
$=\max_{a}\lim_{n\to\infty}|q_{n}(s,a)-q(s,a)|$ (37)
$=\max_{a}|\lim_{n\to\infty}q_{n}(s,a)-q(s,a)|=0$ (38)
The first step holds by substitutions. The inequality is a well known property used to prove that the greedy operator is a contraction, see (Blackwell, 1965). In Equation 37 the limit and max operators can be exchanged because the action space is finite, and finally the limit and absolute value can be exchanged because the absolute value is a continuous function.
C.9 Proof for Theorem 3 for $k=1$ and $\mathcal{I}_{2}$ the identity operator
We will prove Theorem 3, first for $k=1$ for readability, and in the following Appendix C.10 we will extend the proof for $k\geq 1$. In Appendix C.11 we will further extend the proof for value-improvement.
C.9.1 Notation
We use $\mathcal{R}$ to denote the mean-reward vector $\mathcal{R}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, where $\mathcal{R}_{s,a}=\mathbb{E}[R|s,a]$. We use $\mathcal{P}^{\pi}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ to denote the matrix of transition probabilities multiplied by a policy, indexed as follows: $\mathcal{P}^{\pi}_{s,a,s^{\prime},a^{\prime}}=P(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$. We denote the state-action value $q$ and the policy $\pi$ as vectors in the state-action space s.t. $q,\pi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$. The set $\Pi\subset R^{|\mathcal{S}||\mathcal{A}|}$ contains all admissible policies that define a probability distribution over the action space for every state. For convenience, we denote $q(s,a)$ as a specific entry in the vector indexed by $s,a$ and $q(s),\pi(s)$ as the appropriate $|\mathcal{A}|$ dimensional vectors for index $s$. In this notation, we can write expectations as the dot product $q(s)\cdot\pi(s)=\mathbb{E}_{a\sim\pi(s)}[q(s,a)]=v(s)$. With slight abuse of notation, we use $q\cdot\pi=v,\,v\in\mathbb{R}^{|\mathcal{S}|}$ to denote the vector with entries $v(s)$. We use $\max_{a}q\in\mathbb{R}^{|\mathcal{S}|}$ to denote the vector with entries $\max_{a}q(s)=\max_{a}q(s,a)$.
We let $s_{t}$ denote a state $(\cdot,t)\in\mathcal{S}$, that is, a state in the environment arrived at after $t$ transitions. The states $s_{H}$ are terminal states, and the indexing begins from $s_{0}$. We let $q^{m},\pi^{m}$ denote the vectors at iteration $m$ of Algorithm 1. We let $q^{m}_{t},\pi^{m}_{t}$ denote the sub-vectors of all entries in $q^{m},\pi^{m}$ associated with states $s_{t}$. In this notation $q^{1}_{H-1}$ is the $q$ vector for all terminal transitions $(s_{H-1},\cdot)$ after the one iteration of the algorithm.
C.9.2 Proof
Proof.
Convergence for Generalized Policy Iteration with $k=1$
We will prove by backwards induction from the terminal states that the sequence $\lim_{m\to\infty}(\pi^{m},q^{m})$ induced by Algorithm 1 converges for any $q^{0},\pi^{0}$, sufficient greedification operator $\mathcal{I}$ and $k\geq 1$. That is, for every $\epsilon>0$ there exists a $M_{\epsilon}$ such that $\|q^{m}-q^{*}\|\leq\epsilon$ and $\|\pi^{m}\cdot q^{m}-\max_{a}q^{*}\|<\epsilon$ for all $m\geq M_{\epsilon}$, $q^{0}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ and $\pi^{0}\in\Pi$.
Induction Hypothesis: For every $\epsilon>0$ there exist $M^{\epsilon}_{t+1}$ such that for all $m\geq M^{\epsilon}_{t+1}$ we have $\|q_{t+1}^{m}-q_{t+1}^{*}\|\leq\epsilon$, and $\|\pi_{t+1}^{m}\cdot q_{t+1}^{m}-\max_{a}q_{t+1}^{*}\|\leq\epsilon$.
Base Case $t=H-1$: Let $\epsilon>0$. Since states $s_{H}$ are terminal, and have therefore value 0, we have $q_{H-1}^{m}=\mathcal{R}_{H-1}=q_{H-1}^{*}$ and therefore $\|q_{H-1}^{m}-q_{H-1}^{*}\|\leq\epsilon$ trivially holds for all $m\geq 1$.
By the Sufficiency condition of the sufficient greedification operator which induces convergence of $\pi^{m}$ to an $\operatorname*{arg\,max}$ policy with respect to $q$ there exists $M_{H-1}^{\epsilon}$ such that:
$\|\pi_{H-1}^{m}\cdot q_{H-1}^{m}-\max_{a}q_{H-1}^{*}\|=\|\pi_{H-1}^{m}\cdot q_{H-1}^{*}-\max_{a}q_{H-1}^{*}\|\leq\epsilon$
for all $m\geq M_{H-1}^{\epsilon}$. Thus the Induction Hypothesis holds at the base case.
Case $t<H-1$: We will show that if the Induction Hypothesis holds for all states $t+1$, it also holds for states $t$.
Step 1: Let $\epsilon>0$. Assume the Induction Hypothesis holds for states $t+1$. Then there exists $M^{\epsilon}_{t+1}$ such that $\|q^{m}_{t+1}-q^{*}_{t+1}\|\leq{\epsilon}$ and $\|\pi_{t+1}^{m}\cdot q_{t+1}^{m}-\max_{a}q_{t+1}^{*}\|\leq\epsilon$ for all $m\geq M^{\epsilon}_{t+1}$.
Let us define the transition matrix $\mathcal{P}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}|}$ with $\mathcal{P}_{s,a,s^{\prime}}=P(s^{\prime}|s,a)$.
First, for all $m\geq M^{\epsilon}_{t+1}$ we have:
$\|q_{t}^{{m+1}}-q^{*}_{t}\|$ $=\|\mathcal{R}+\gamma\mathcal{P}(\pi_{t+1}^{m+1}\cdot q_{t+1}^{m})-\mathcal{R}-\gamma\mathcal{P}\max_{a}q^{*}_{t+1}\|$ (39)
$=\gamma\|\mathcal{P}(\pi_{t+1}^{m+1}\cdot q_{t+1}^{m})-\mathcal{P}\max_{a}q^{*}_{t+1}\|$ (40)
$\leq\|\mathcal{P}\|\|\pi_{t+1}^{m+1}\cdot q_{t+1}^{m}-\max_{a}q^{*}_{t+1}\|$ (41)
$\leq\|\pi_{t+1}^{m+1}\cdot q_{t+1}^{m}-\max_{a}q^{*}_{t+1}\|$ (42)
$\leq\epsilon$ (43)
(39) is by substitution based on step 4 in Algorithm 1 for $k=1$. (41) is by the definition of the operator norm $\|\mathcal{P}\|$. (42) is by the fact that the operator norm in sup-norm of all transition matrices is 1 (Bertsekas, 2007). (43) is slightly more involved, and follows from the Induction Hypothesis and the limit-sufficient greedification.
Let us show that (43), i.e. $\|\pi^{m+1}_{t+1}\cdot q^{m}_{t+1}-\max_{a}q^{*}_{t+1}\|\leq\epsilon$ holds. Under the infinity norm holds point-wise for each state $s\in\mathcal{S}$:
$-\epsilon$ $\leq[\pi^{m}_{t+1}\cdot q^{m}_{t+1}](s)-\max_{a}q^{*}_{t+1}(s,a)$ (44)
$\leq\bm{[\pi^{m+1}_{t+1}\cdot q^{m}_{t+1}](s)-\max_{a}q^{*}_{t+1}(s,a)}$ (45)
$\leq\max_{a^{\prime}}q^{m}_{t+1}(s,a^{\prime})-\max_{a}q^{*}_{t+1}(s,a)$ (46)
$\leq\max_{a^{\prime}}\big{(}q^{m}_{t+1}(s,a^{\prime})-q^{*}_{t+1}(s,a^{\prime})\big{)}$ (47)
$\leq\epsilon\,.$ (48)
(44) is the induction hypothesis $\|\pi^{m}_{t+1}\cdot q^{m}_{t+1}-\max_{a}q^{*}_{t+1}\|\leq\epsilon$, which holds under the infinity norm point wise, (45) uses the sufficient greedification operatorproperty $[\pi^{m}_{t+1}\cdot q^{m}_{t+1}](s)\leq[\pi^{m+1}_{t+1}\cdot q^{m}_{t+1}](s)$, (46) the inequality $[\pi^{m+1}_{t+1}\cdot q^{m}_{t+1}](s)\leq\max_{a^{\prime}}q^{m}_{t+1}(s,a^{\prime})$, (47) the inequality $-\max_{a}q^{*}_{t+1}(s,a)\leq-q^{*}_{t+1}(s,a^{\prime}),\forall a^{\prime}\in\mathcal{A}$, and (48) the induction hypothesis $\|q^{*}_{t+1}-q^{m}_{t+1}\|\leq\epsilon$.
Step 2: Pick $M_{t}^{\epsilon}\geq M^{\epsilon}_{t+1}$ such that for all $m\geq M_{t}^{\epsilon}$ we have $\|\pi_{t}^{m}\cdot q_{t}^{m}-\max_{a}q^{*}_{t}\|\leq\epsilon$ which must exist by limit-sufficiency’s condition, Step 1 and the Induction Hypothesis. Thus, the Induction Hypothesis holds for all states $t$ if it holds for states $t+1$.
Finally, let $\epsilon>0$. By backwards induction, for each $t=0,\dots,H-1$ there exists $M_{\epsilon}^{t}$ such that for all $m\geq M_{\epsilon}^{t}$ we have $\|q_{t}^{m}-q_{t}^{*}\|\leq\epsilon$, and $\|\pi_{t}^{m}\cdot q_{t}^{m}-\max_{a}q_{t}^{*}\|\leq\epsilon$. Therefore, we can pick $N_{\epsilon}=\max_{t=0,\dots,H-1}M_{\epsilon}^{t}$ such that $\|q_{t}^{m}-q_{t}^{*}\|\leq\epsilon$, and $\|\pi_{t}^{m}\cdot q_{t}^{m}-\max_{a}q_{t}^{*}\|\leq\epsilon$ for all $m\geq N_{\epsilon}$ and $t=0,\dots,H-1$, proving that Algorithm 1 converges to an optimal policy and optimal q-values for any $\pi^{0}\in\Pi,q^{0}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|},k=1$ and sufficient greedification $\mathcal{I}$. ∎
We proceed to extend the proof for $k\geq 1$ below.
C.10 Extension of the Proof for Theorem 3 to $k\geq 1$ and $\mathcal{I}_{2}$ the identity operator
In this section we will extend the proof of Theorem 3 from Appendix C.9 to $k\geq 1$. Much of the proof need not be modified. In order to extend the proof to $k\geq 1$, we only need to show the following: For all $k\geq 1$ and every $\epsilon>0$ such that the Induction Hypothesis holds, there exists an $M_{\epsilon}^{t}$ such that $\|q_{t}^{{m+1}}-q^{*}_{t}\|\leq\epsilon$.
Proof.
We will first extend the notation: let $q^{m,i}_{t}$ denote the vector $q$ at states $t$ after $m$ algorithm iterations and $i\geq 1$ Bellman updates, such that $q^{m,i}_{t}=(\mathcal{T}^{\pi_{1}}q^{m,i-1})_{t},\,q^{m,0}_{t}=q^{m}_{t}$ and finally $q^{m+1}_{t}=q^{m,k}_{t}$.
Second, we will extend the Induction Hypothesis:
Extended Induction Hypothesis: For every $\epsilon>0$ there exist $M^{\epsilon}_{t+1}$ such that for all $m\geq M^{\epsilon}_{t+1}$ and $i\geq 0$ we have $\|q_{t+1}^{m,i}-q_{t+1}^{*}\|\leq\epsilon$, and $\|\pi_{t+1}^{m}\cdot q_{t+1}^{m,i}-\max_{a}q_{t+1}^{*}\|\leq\epsilon$.
The Base Case does not change, so we will proceed to Step 1 in the Inductive Step. We need to show that there exists an $M_{t}^{\epsilon}$ such that $\|q^{m,i}_{t}-q^{*}_{t}\|\leq\epsilon$ for all $i\geq 0$ and $m\geq M_{t}^{\epsilon}$.
Let $\epsilon>0$ and $m\geq M_{t}^{\epsilon}\geq M_{t+1}^{\epsilon}$.
First, for any $i\geq 1$:
$\|q_{t}^{{m,i}}-q^{*}_{t}\|$ $=\|\mathcal{R}+\gamma\mathcal{P}(\pi_{t+1}^{m+1}\cdot q_{t+1}^{m,i-1})-q^{*}_{t}\|$
$\leq\|\mathcal{P}\|\|\pi_{t+1}^{m+1}\cdot q_{t+1}^{m,i-1}-\max_{a}q^{*}_{t+1}\|$
$\leq\epsilon$
The first equality is the application of the Bellman Operator in line 4 in Algorithm 1 the $i$th time. The rest follows from Proof C.9 and the extended Induction Hypothesis.
Second, we need to show that this holds for $i=0$ as well:
$\|q_{t}^{{m,0}}-q^{*}_{t}\|=\|q_{t}^{{m-1,k}}-q^{*}_{t}\|\leq\|\pi_{t+1}^{m}\cdot q_{t+1}^{m-1,k-1}-\max_{a}q^{*}_{t+1}\|\leq\epsilon$
The first equality is by definition, and the the first and second inequalities are by the same argumentation as above.
The rest of the proof need not be modified. ∎
C.11 Extension for $\mathcal{I}_{2}$ a general improvement operator
We extend the proof from the above section for all non-detriment operators (that is, non-strict greedification operators) $\mathcal{I}_{2}$ used for value improvement.
Proof.
Similarly to the proof of Theorem 3 from Appendix C.9 (and C.10) we will prove by backwards induction from the terminal states $s_{H}$ that the sequence $\lim_{m\to\infty}(\pi^{m},q^{m})$ induced by Algorithm 2 converges for any $q^{0},\pi^{0}$, sufficient greedification operator $\mathcal{I}_{1}$, greedification operator $\mathcal{I}_{2}$ and $k\geq 1$. That is, for every $\epsilon>0$ there exists a $M_{\epsilon}$ such that $\|q^{m}-q^{*}\|\leq\epsilon$ and $\|\pi^{m}\cdot q^{m}-\max_{a}q^{*}\|<\epsilon$ for all $m\geq M_{\epsilon}$, $q^{0}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ and $\pi^{0}\in\Pi$. The proof follows directly from the proof in Appendix C.9. The base case is not modified - the $q$s converge immediately and the policy convergence is not influenced by the introduction of $\mathcal{I}_{2}$. The Induction Hypothesis need not be modified. In the inductive step, Step 1 follows directly from the Induction Hypothesis, and Step 2 need not be modified for the same reason the base case need not be modified. ∎
C.12 Convergence of Algorithm 2 with Bounded Greedification operators
We extend the proof from Appendices C.9 and C.10 to bounded greedification operators.
C.12.1 Bounded Greedification converges to an $\operatorname*{arg\,max}$ policy in finite steps
We will begin by proving that operators with the Bounded Greedification property:
$|\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)-\sum_{a\in\mathcal{A}}\pi(a|s)q(s,a)|>\epsilon,$
unless $\sum_{a\in\mathcal{A}}\mathcal{I}(\pi,q)(a|s)q(s,a)=\max_{a}q(s,a)$ are guaranteed to convergence to an $\operatorname*{arg\,max}$ policy with respect to any $q\in\mathcal{Q}$, in a finite number of steps.
Lemma 8.
Let $\mathcal{I}$ be a bounded greedification operator and let a sequence $\pi_{n+1}=\mathcal{I}(q,\pi_{n})$. For any starting $\pi_{0}\in\Pi,q\in\mathcal{Q}$, there exists an $M$ for which:
$\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)=\max_{a}q(s,a),\quad\forall n>M.$
Proof.
Let $\mathcal{I}$ be a Bounded Greedification operator. At each iteration, the sequence $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)$ must increase, i.e. $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)>\sum_{a\in\mathcal{A}}\pi_{n-1}(a|s)q(s,a),n>0$, for at least one state $s\in\mathcal{S}$. The same sequence is monotonically non-decreasing, by definition of greedification, for all other states. Therefore, the sequence $\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)$ is monotonically increasing (for each state $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)$ is at least as large as in the past step, and in at least one state it is distinctly higher), unless $\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)=\max_{a}q(s,a)$.
Due to the Bounded Greedification property, the minimum increase is bounded by $\epsilon$, that is:
$\min_{\pi_{n}}|\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)-\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\pi_{n-1}(a|s)q(s,a)|>\epsilon,\quad n>0.$
The sequence $\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\pi_{n}(a|s)q(s,a)$ is bounded by $\sum_{s\in\mathcal{S}}\max_{a}q(s,a)$ from above, and by $\sum_{s\in\mathcal{S}}\min_{a}q(s,a)$ from below.
Since the sequence is bounded from below and above and increases by a bounded amount $\epsilon>0$, it must converge in a finite $n<\infty$. ∎
C.12.2 Modified Induction for Bounded Greedification
We modify the induction of the proof of Theorem 3 with finite-sufficient greedification operators, that converge to an $\operatorname*{arg\,max}$ policy in a finite number of iterations.
Proof.
Modified Induction Hypothesis: There exist $M_{t+1}$ such that for all $m\geq M_{t+1}$ we have $q_{t+1}^{m}=q_{t+1}^{*}$, and $\pi_{t+1}^{m}\cdot q_{t+1}^{m}=\max_{a}q_{t+1}^{*}$.
Modified Base Case: Because the convergence to the $\operatorname*{arg\,max}$ is in finite time (Lemma 8) there exists $M_{H-1}$ such that:
$\pi_{H-1}^{m}\cdot q_{H-1}^{m}=\max_{a}q_{H-1}^{*}$
for all $m\geq M_{H-1}$. Thus the Modified Induction Hypothesis holds at the base case.
Modified Case $t<H-1$ Step (1): Similarly, for all $m\geq M_{t+1}$ we have:
$\|q_{t}^{{m+1}}-q^{*}_{t}\|$ $=\|\mathcal{R}+\gamma\mathcal{P}(\pi_{t+1}^{m+1}\cdot q_{t+1}^{m})-\mathcal{R}-\gamma\mathcal{P}\max_{a}q^{*}_{t+1}\|$ (49)
$=\gamma\|\mathcal{P}(\pi_{t+1}^{m+1}\cdot q_{t+1}^{m})-\mathcal{P}\max_{a}q^{*}_{t+1}\|$ (50)
$\leq\|\mathcal{P}\|\|\pi_{t+1}^{m+1}\cdot q_{t+1}^{m}-\max_{a}q^{*}_{t+1}\|$ (51)
$=0$ (52)
Since also $\|q_{t}^{{m+1}}-q^{*}_{t}\|\geq 0$, we have $\|q_{t}^{{m+1}}-q^{*}_{t}\|=0$ and $q_{t}^{{m+1}}=q^{*}_{t}$.
Step (2): Pick $M_{t}\geq M_{t+1}$ such that for all $m\geq M_{t}$ we have $\|\pi_{t}^{m}\cdot q_{t}^{m}-\max_{a}q^{*}_{t}\|=0$ which must exist due to convergence to the argmax in finite time of this operator class. Thus, the Modified Induction Hypothesis holds for all states $t$ if it holds for states $t+1$. ∎
Appendix D Experimental Details
D.1 Gradient-Based VI-TD3
Gradient-based VI-TD3 copies the existing policy used to compute value targets (the target policy, in TD3) $\pi_{\theta^{\prime}}$ into a new policy $\pi_{\theta^{\prime}}^{\prime}$. The algorithm executes N repeating gradient steps on $\pi_{\theta^{\prime}}^{\prime}$ with respect to states $s_{t+1}\in b$ with the same operator TD3 uses to improve the policy (the deterministic policy gradient) and with respect to the same batch $b$. The value-improved target $y(s_{t},a_{t})$ is computed in the same manner to the original target of TD3 but with the fresh greedified target network $\pi_{\theta^{\prime}}^{\prime}$. In TD3, that summarizes to sampling an action from a clipped Gaussian distribution with mean $\pi_{\theta^{\prime}}^{\prime}(s_{t+1})$, variance parameter $\sigma$ and clipped between $(-\beta,\beta)$:
$a^{\prime}\sim\mathcal{N}(\pi_{\theta^{\prime}}^{\prime}(s_{t+1}),\sigma).clip(-\beta,\beta)$ (53)
And using the action $a^{\prime}$ to compute the value target in the Sarsa manner:
$y(s_{t},a_{t})=r_{t}+\gamma\min_{i\in\{1,2\}}q_{\phi_{i}}(s_{t+1},a^{\prime}),\forall(s_{t},a_{t},r_{t},s_{t+1})\in b$ (54)
The policy used to compute the value targets $\pi_{\theta^{\prime}}$ is then discarded.
D.2 Sample-based $\operatorname*{arg\,max}$
The sampling based $\operatorname*{arg\,max}$ (approximate) greedification operator acts as follows: First, sample $N$ actions from the evaluation policy $a_{1},\dots,a_{N}\sim\pi$. In TD3, we use the same policy used to compute value targets $\mathcal{N}(\pi_{\theta^{\prime}}(s_{t+1}),\sigma).clip(-\beta,\beta)$, see Appendix D.1. Second, find the action with highest $q$ value: $a_{max}=\operatorname*{arg\,max}_{i}\min_{\phi_{j}}q_{\phi_{j}}(s_{t+1},a_{i})$. Finally, the improved policy used to compute the improved targets is $\mathcal{N}(a_{max},\sigma).clip(-\beta,\beta)$, in the manner of TD3. In our experiments, $N=128$ samples were used.
D.3 Implicit Policy Improvement with Expectile Loss
The expectile-loss $\mathcal{L}^{\tau}_{2}$ proposed by Kostrikov et al. (2022) as an implicit policy improvement operator for continuous-domain Q-learning can be formulated as follows: when $y(s_{t},a_{t})>q(s_{t},a_{t})$ (the target is greater than the prediction), the loss equals $\tau(y(s_{t},a_{t})-q(s_{t},a_{t}))^{2}$. When $y(s_{t},a_{t})<q(s_{t},a_{t})$ (the target is smaller than the prediction) the loss equals $(1-\tau)(y(s_{t},a_{t})-q(s_{t},a_{t}))^{2}$. If $\tau=0.5$, this loss is equivalent to the baseline $\mathcal{L}_{2}$ loss. Intuitively, when $\tau>0.5$ the agent favors errors where the prediction should increase, over predictions where it should reduce. I.e. the agent favors targets where $\pi^{\prime}(s_{t+1})$ (the implicit policy evaluated on the next state) chooses "better" actions than the current policy, directly approximating the value of an improved policy.
By imposing this loss on the value network, in stochastic environments the network may learn to be risk-seeking, by implicitly favoring interactions $s_{t},a_{t},r_{t},s_{t+1}$ where the observed $r_{t}$ was large or the state $s_{t+1}$ was favorable. This is addressed by Kostrikov et al. (2022) by learning an additional $v_{\psi}$ network that is trained with the expectile loss, while the $q$ network is trained with SARSA targets $r_{t}+\gamma v_{\psi}(s_{t+1})$ and the regular $\mathcal{L}_{2}$ loss, while the $v_{\psi}$ network is trained with targets $y(s_{t},a_{t})=q_{\phi}(s_{t},a_{t})$ and the expectile loss. In deterministic environments this is not necessary however, and in our experiments we have directly replaced the $\mathcal{L}_{2}$ loss on the value $q_{\phi}$ with the expectile loss.
The value target $y(s_{t},a_{t})$ remained the unmodified target used by TD3 / SAC respectively.
D.4 Evaluation Method
We plot the mean and standard error for evaluation curves across multiple seeds. Evaluation curves are computed as follows: after every $n=5000$ interactions with the environment, $m=3$ evaluation episodes are ran with the latest network of the agent (actor and critic). The score of the agent is the return averaged across the $m$ episodes. The actions in evaluation are chosen deterministically for TD3, SAC and TD7 with the mean of the policy (the agents use Gaussian policies). The evaluation episodes are not included in the agent’s replay buffer or used for training, nor do they count towards the number of interactions.
D.5 Implementation & Hyperparemeter Tuning
Our implementation of TD3 and SAC relies on the popular code base CleanRL (Huang et al., 2022). CleanRL consists of implementations of many popular RL algorithms which are carefully tuned to match or improve upon the performance reported in the original paper. The implementations of TD3 and SAC use the same hyperparameters as used by the authors (Fujimoto et al. (2018) and Haarnoja et al. (2018a) respectively), with the exception of the different learning rates for the actor and the critic in SAC, which were tuned by CleanRL.
For the TD7 agent, we use the original implementation by the authors (Fujimoto et al., 2023), adapting the action space to the DeepMind control’s in the same manner as CleanRL’s implementation of TD3. Additionally, a non-prioritized replay buffer has been used for TD7 which was used by the TD3 and SAC agents as well. The hyperparameters are the same as used by the author.
The VI-variations of all algorithms use the same hyperparameters as the baseline algorithms without any additional tuning, with the exception of grid search for the greedification parameters $\tau$ presented in Figure 3.
D.6 Network Architectures
The experiments presented in this paper rely on standard architectures for every baseline. TD3 and SAC used the same architecture, with the exception that SAC’s policy network predicts a mean of a Gaussian distribution as well as standard deviation, while TD3 predicts only the mean. TD7 used the same architecture proposed and used by Fujimoto et al. (2023).
TD3 and SAC:
Actor: 3 layer MLP of width 256 per layer, with ReLU activations on the hidden layers. The final action prediction is passed through a $\tanh$ function.
Critic: 3 layer MLP of width 256 per layer, with ReLU activations on the hidden layers and no activation on the output layer.
TD7: Has a more complex architecture, which is specified in (Fujimoto et al., 2023).
D.7 Hyperparemeters
TD3 SAC TD7 exploration noise $0.1$ exploration noise $0.1$ Target policy noise $0.2$ Target policy noise $0.2$ Target smoothing $0.005$ Target smoothing $0.005$ noise clip $0.5$ auto tuning of entropy True noise clip $0.5$ Critic learning rate 1e-3 Critic learning rate 3e-4 Learning rate 3e-4 Policy learning rate 3e-4 Policy learning rate 3e-4 Policy update frequency 2 Policy update frequency 2 Policy update frequency 2 $\gamma$ $0.99$ $\gamma$ $0.99$ $\gamma$ $0.99$ Buffer size $10^{6}$ Buffer size $10^{6}$ Buffer size $10^{6}$ Batch size 256 Batch size 256 Batch size 256 learning start $10^{4}$ learning start $10^{4}$ learning start $10^{4}$ evaluation frequency $5000$ evaluation frequency $5000$ evaluation frequency $5000$ Num. eval. episodes $3$ Num. eval. episodes $3$ Num. eval. episodes $3$